# Entropy-guided $k$-Guard Sampling for Long-Horizon Autoregressive Video Generation

### v1

@@ 2.1 Video World Models
Video-based world models aim to learn an internal representation of an environment, allowing the system to predict future states, simulate interactions, and support planning (Ha and Schmidhuber, 2018; Ding et al., 2024; Long et al., 2025; Zhang et al., 2025). Recent progress in large-scale video generation has enabled the creation of high-fidelity world simulators capable of producing visually realistic and physically plausible sequences (OpenAI, 2024), which are particularly valuable for applications such as autonomous driving and robotics (wang2024worlddreamer; Li et al., 2025).
A wide range of generative architectures have been explored for video synthesis. Diffusion models and Generative Adversarial Networks (GANs) have shown success in producing high-quality frame sequences (Ho et al., 2022; Clark and Fidler, 2019), while autoregressive (AR) models have emerged as a powerful alternative due to their inherent capacity to model temporal dependencies. By factorizing the joint distribution of video tokens or frames into a product of conditional probabilities, AR models generate coherent, long-duration videos either frame-by-frame (Gu et al., 2025) or token-by-token (Wu et al., 2024), enabling fine-grained temporal control and interactive generation.
@@ A.1 Experimental Protocols (Concise)
A.2 Use of LLMs (Disclosure)
LLMs were used for language polishing and minor code refactoring suggestions only; authors verified all outputs. No dataset labeling or claim-critical content was delegated without human validation.
A.3 Ethics (Brief)
We comply with dataset licenses and anonymization policies; no personally identifiable information was used. Potential risks (e.g., misuse in safety-critical scenarios) are discussed and bounded by research-only usage.
@@ C.1 General-Domain Models
To validate the effectiveness of ENkG in general domains, we evaluate it on Lumos-1 (Yuan et al., 2025) and NBP (Ren et al., 2025a). Our method mitigates error accumulation during long-horizon video generation and produces more temporally consistent results with reduced color drift and collapse artifacts. Representative qualitative results are shown in Figure 9 and Figure 10.
@@ D.2 Consequences of the Low-Entropy Trap
In video generation, the low-entropy trap does not merely lead to trivially frozen frames. More subtly, it manifests as: (i) background smearing, where large regions of the background collapse into blurry blobs and lose meaningful structural detail; (ii) global color shift, where the entire frame drifts toward an unnatural color cast, making consecutive frames look consistently tinted or over-/under-exposed; and (iii) texture freezing, where fine-grained appearance patterns (e.g., grass, water, sky) become unnaturally static and appear glued to the camera or object rather than evolving with the motion. These effects are illustrated in Figure 12 and 14, where the model settles into visually coherent yet clearly undesirable trajectories.

### v2

@@ 2.1 Video World Models
Video-based world models aim to learn an internal representation of an environment, allowing the system to predict future states, simulate interactions, and support planning (Ha and Schmidhuber, 2018; Ding et al., 2024; Long et al., 2025; Zhang et al., 2025). Recent progress in large-scale video generation has enabled the creation of high-fidelity world simulators capable of producing visually realistic and physically plausible sequences (OpenAI, 2024), which are particularly valuable for applications such as autonomous driving and robotics (Li et al., 2025).
A wide range of generative architectures have been explored for video synthesis. Diffusion models and Generative Adversarial Networks (GANs) have shown success in producing high-quality frame sequences (Ho et al., 2022), while autoregressive (AR) models have emerged as a powerful alternative due to their inherent capacity to model temporal dependencies. By factorizing the joint distribution of video tokens or frames into a product of conditional probabilities, AR models generate coherent, long-duration videos either frame-by-frame (Gu et al., 2025) or token-by-token (Wu et al., 2024), enabling fine-grained temporal control and interactive generation.
@@ C.1 General-Domain Models
To validate the effectiveness of ENkG in general domains, we evaluate it on Lumos-1 (Yuan et al., 2025) and NBP (Ren et al., 2025a). Our method mitigates error accumulation during long-horizon video generation and produces more temporally consistent results with reduced color drift and collapse artifacts. Representative qualitative results are shown in Figure 9.
@@ D.2 Consequences of the Low-Entropy Trap
In video generation, the low-entropy trap does not merely lead to trivially frozen frames. More subtly, it manifests as: (i) background smearing, where large regions of the background collapse into blurry blobs and lose meaningful structural detail; (ii) global color shift, where the entire frame drifts toward an unnatural color cast, making consecutive frames look consistently tinted or over-/under-exposed; and (iii) texture freezing, where fine-grained appearance patterns (e.g., grass, water, sky) become unnaturally static and appear glued to the camera or object rather than evolving with the motion. These effects are illustrated in Figure 11 and 13, where the model settles into visually coherent yet clearly undesirable trajectories.